please help solve this
3x=wyz
solve for x, w, and y
please explain so i can try to do others without help, thank you very much

Answer by kietra(57)

3x=wyz
solve for x, w, and y
To solve for a variable, say x, you are going to work the problem around so you have x= something. You won't be able to get rid of all the letters in this case.
First, solve for x.
To get x by itself, you'll have to divide by 3 on both sides.
3x = wyz
x = wyz/3
To solve for w, divide yz on both sides
3x = wyz
3x/yz = w
To solve for y, divide both sides by wz.
3x = wyz
3x/wz = y
If you have more questions, come back!